How many 4-digit numbers can be formed from the digits 0, 1, 2, 3, 4, 5 if the first digit must not be 0 and repetition of digits is not allowed?
a.300
b.320
c.360
d.280
1
Expert's answer
2020-12-08T06:36:22-0500

Question: How many 4-digit numbers can be formed from the digits 0, 1, 2, 3, 4, 5 if the first digit must not be 0 and repetition of digits is not allowed? 


I.

In place of the first digit in the number, we can place 5 digits from the list, since the first digit cannot be 0. We use the placement rule without repetitions.

"A_5^1=n!\/(n-m)!=5!\/(5-1)!=5"

For the next ones, we do the same, only we take into account the fact that one digit has gone into the first cell, and there are 5 available. We also use the formula placement 3 by 5.

"\u0410_5^3=5!\/(5-3)!=3*4*5=60"

Multiply two counted numbers and get the number of possible numbers with a given set of digits.

"A=5*60=300"

Answer: a. 300


II. Another method.

Let's consider each digit of a 4-digit number separately:

 "a_1\not=[0]" and can take the value "[1,2,3,4,5] \implies a_1=5"

"a_2\not=a_1" and can take the value "[0,1,2,3,4,5]" , but "a_1" is equal to one digit, and under our condition the numbers should not be repeated. "\implies a_2=5"

"a_3\not=a_1\not= a_2" the same as with "a_2" , only one less than "a_3=4"

"a_4\not=a_1\not=a_2\not= a_3" in the same way as with the rest "\implies a_4=3"

Now we multiply all the values of ai.

"a_1*a_2*a_3*a_4=5*5*4*3=300"

Answer: a. 300